OpenStudy (anonymous):

A researcher wants to conduct a genetic study using 25 randomly-selected volunteers. He has a volunteer pool of 30, composed of 15 males and 15 females. What is the probability of 12 male and 13 female volunteers OR 13 male and 12 female volunteers being selected at random for the study? A. 17/522 B. 455/522 C. 105/522 D. 175/261

OpenStudy (jim766):

do you know about combinations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@xGuardians

OpenStudy (anonymous):

@iambatman

OpenStudy (jim766):

15C12 * 15C13 + 15C13 * 15C12 ________________ ________________ 30C 25 30 C 25

OpenStudy (kropot72):

This problem is sampling without replacement. \[P(12\ males)=\frac{15C12\times15C13}{30C25}=\frac{\frac{15!}{12!3!}\times\frac{15!}{13!2!}}{\frac{30!}{25!5!}}\] The probability of 13 males is the same as for 12 males, therefore the required probability is found by doubling the result for 12 males.

OpenStudy (anonymous):

@ajaylounsber

OpenStudy (anonymous):

yeh?

OpenStudy (anonymous):

can't help this time sorry bro

OpenStudy (kropot72):

The calculation for the probability of 12 male and 13 female volunteers OR 13 male and 12 female volunteers being selected at random for the study is: \[2\times\frac{15\times14\times13\times15\times14\times5\times4\times3\times2}{3\times2\times2\times30\times29\times28\times27\times26}=you\ can\ calculate\]